# Tunable Molecular Interactions Near an Atomic Feshbach Resonance: Stability and Collapse of a Molecular Bose-Einstein Condensate

### v1

@@ Abstract
Understanding and controlling interactions of ultracold molecules has been a central goal in quantum chemistry research. Recent experiments on atoms near a Feshbach resonance offer the key to prepare and investigate molecules in the quantum many-body regime. Just as Feshbach resonances allow tuning of the scattering length of bosonic atoms, we show that they also modify the scattering length of Feshbach molecules which are constituted from these atoms. Based on calculations of the compressibility, we determine the stability phase diagrams of molecular condensates and show that their instability can be associated with a sign change of the inter-molecular interactions. We derive universal expressions for the molecular scattering lengths, presented in terms of experimentally measurable quantities. These will enable control of interactions between Feshbach molecules as well as further studies of few- and many-body reactions involving Feshbach molecules in the quantum regime.
Introduction — Molecular condensates formed from bosonic atoms present exciting opportunities for revealing new states of matter and novel quantum phase transitions Radzihovsky et al. (2004); Romans et al. (2004); Radzihovsky et al. (2008); Radzihovsky and Choi (2009). A key technique for producing these diatomic molecules is through Feshbach resonances Chin et al. (2010), where magnetoassociation converts pairs of ultracold atoms into bound diatomic molecules. This process, involving a sweep of a magnetic field across a resonance, has enabled the preparation of stable molecular condensates in equilibrium Zhang et al. (2021, 2023). Such condensates in more direct proximity to resonance are inaccessible due to severe three-body recombination Fedichev et al. (1996); Bedaque et al. (2000); Weber et al. (2003); Fletcher et al. (2013) and associated particle loss and heating. This situation is in striking contrast to the two-component Fermi gas, where, due to the Pauli exclusion, stable molecular condensates  Petrov et al. (2004, 2005), and as a result the associated BCS-BEC crossover, have been extensively studied Zwerger (2011); Regal et al. (2004); Zwierlein et al. (2004); Chen et al. (2005).
The success reported in Refs. Zhang et al. (2021, 2023) then raises the natural question of how to understand and characterize the stability or, more typically, lack thereof, in these near-resonant molecular condensates  Jeon et al. (2002); Koetsier et al. (2009). In this paper we address this important issue, focusing on the instability of molecular condensates as represented by the behavior of the molecule-molecule scattering length. This is inspired by the behavior of atomic condensates, whose two-body $s$-wave scattering length $a_{s}$ can generally be expressed as
$a_{s}$ $=a_{\mathrm{bg}}\big{(}1-\frac{\Delta B}{B-B_{0}}\big{)},$ (1)
where $B_{0}$ is the resonance value of the magnetic field $B$, $\Delta B$ is the resonance width and $a_{\mathrm{bg}}$ is the atomic $s$-wave background scattering length. The leading-order contribution to $a_{s}$ due to the Feshbach coupling can be represented by Fig. 1(a)  [^1] Higher-order contributions involve multiple repetitions of the process in Fig. 1(a). While these contributions do not alter the resonant structure of $a_{s}$ in Eq. (1),they do lead to a renormalization of $B_{0}$. . When $a_{s}$ is negative, the atomic condensates undergo a collapse Kagan et al. (1998), which occurs when the magnetic field $B$ is above the resonance $B_{0}$ by no more than the resonance width $\Delta B$, or $B_{0}<B<B_{0}+\Delta B$ (for a positive atom background scattering length $a_{\mathrm{bg}}>0$). For concreteness, here we have assumed $\Delta B>0$ so that the energy of the bare molecules falls below that of the atom continuum when the magnetic field is below the resonance $B<B_{0}$.   [^2] The formulae derived here can be easily generalized to the case of negative $\Delta B$.. What has not been widely appreciated is that the Feshbach coupling also resonantly modifies the scattering properties of molecules and can lead to an instability of a molecular Bose-Einstein condensate.
@@
Figure 1: Scattering between atoms (a) and molecules (b) near an atomic Feshbach resonance. (a) Leading-order scattering process that contributes to the resonant term appearing in $a_{s}$ in Eq. (1). This involves two incoming atoms (single blue circles) that temporarily combine into one molecule (two attached circles) through the Feshbach coupling and then dissociate into two outgoing free atoms. The orange region represents virtual processes. (b) The corresponding leading-order contribution to the molecular scattering length $a_{\mathrm{mm}}$ near the resonance. Here, two molecules approach each other, temporarily break up into four free atoms which propagate, interact, and then recombine into another two molecules. As in (a), this process universally depends on the Feshbach coupling, but is of higher order in the coupling strength.
In this paper, on the basis of a two-channel model, we demonstrate that, similar to the atomic condensate case, Feshbach resonances also alter the inter-molecular scattering length $a_{\mathrm{mm}}$. This can lead to the instability of molecular condensates when the molecular scattering length is tuned to a negative value. Importantly, this molecular scattering length can be further modified due to finite density effects which we characterize through an additional effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$. This effective scattering length will depend on the density significantly for a narrow resonance due to many-body effects. In the case of a very wide Feshbach resonance this dependence is weak and can be safely ignored, so that $a_{\mathrm{mm}}^{\mathrm{eff}}$ becomes equivalent to the two-body molecular scattering length $a_{\mathrm{mm}}$.
Arriving at the scattering lengths $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{\mathrm{mm}}$ presents a significant challenge for theory. Hints of this difficulty can already be seen from Fig. 1(b), which indicates just one of the leading order processes which is already of higher order in the coupling strength than the simpler atomic analogue in Fig. 1(a). Additional challenges come from the need to include contributions from background interactions among atoms and molecules.
Summary of main results — Our calculations are based on a variational treatment, from which we address the behavior of the compressibility and deduce the associated scattering lengths. We first summarize some general results.
Figure 2: Phase diagrams and molecular scattering lengths. (a,b) Ground-state stability phase diagrams for (a) wide and (b) narrow resonances. Plotted in (a,b) is a map of the compressibility $\kappa=dn/d\mu$ as a function of atom number density $n$ and the detuning $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$. The compressibility $\kappa$ and the detuning $\bar{\nu}$ are normalized by $\kappa_{\mathrm{bg}}=m_{1}/(4\pi\hbar^{2}a_{\mathrm{bg}})$ and the resonance width $\Delta_{\bar{\nu}}=\Delta\mu_{m}\Delta B$, respectively. The atomic condensate is nonzero only in the ASF phase, while the molecule condensate is nonzero in both the ASF and MSF phases. Stable regions with positive compressibility $\kappa>0$ are shaded in orange, while the unstable region is in blue. The red dashed lines in (a,b) correspond to the quantum critical point at $\bar{\nu}_{c}(n)$, separating the ASF from the MSF phase, and are determined independently of stability. The parameters $\bar{\nu}_{c,-}$ and $\bar{\nu}_{c,+}$ represent the detuning separating the unstable and stable regions. (c,d) Corresponding two-body scattering length $a_{\mathrm{mm}}$, and its many-body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ at a typical density $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$. For the narrow resonances in (b,d) we choose parameters from the resonance of 133Cs at $B_{0}=19.849$ G used in Refs. Zhang et al. (2021, 2023); Wang et al. (2024), while for the wide resonances in (a,c) we increased the resonance width $\Delta_{\bar{\nu}}$ from its value in (b,d) by $10^{2}$ times, keeping all other parameters the same.
@@
Plotted in Fig. 2(a,b) are the zero-temperature stability phase diagrams for both a wide and narrow resonance, in terms of density $n$ on the vertical axis and detuning $\bar{\nu}$ on the horizontal axis. These plots are obtained by numerical calculations of the compressibility $\kappa$. Both panels indicate a stable molecular superfluid (MSF) phase at large negative detuning, a stable atomic superfluid (ASF) phase at large positive detuning, and an unstable region in between, near the resonance. The instability corresponds to parameters where the compressibility is negative $\kappa<0$. We call attention to the detuning labels $\bar{\nu}_{c,+}$ and $\bar{\nu}_{c,-}$ which delineate the two end points for the unstable region. In Fig. 2(c,d), we plot both our calculated two-body scattering length, $a_{\mathrm{mm}}$, and its many body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$.
@@
There is a pronounced similarity between panels (a) and (b) of Fig. 2 on the atomic side, as for both narrow and wide resonances, the phase boundary of the unstable region is essentially governed by the sign change of the two-body atomic scattering length $a_{s}$. Here, in both cases the upper boundary detuning $\bar{\nu}_{c,+}$ of the unstable regime is, thus, nearly $n$-independent.
@@
By contrast, the behavior of the compressibility on the molecular side of the phase diagrams in Fig. 2 displays a notable contrast between the narrow and wide resonance cases. We summarize some of our key observations as follows:
  1. 1.
@@
    In the wide resonance case the entire region of instability shows very little density dependence. For this reason we should be able to characterize the molecular-condensate stability entirely in terms of a two body scattering length (that is appropriate to zero density), as is similar to the atomic condensate case. This can be seen from Fig. 2(c) which shows the near-equivalence $a_{\mathrm{mm}}^{\mathrm{eff}}(n)\approx a_{\mathrm{mm}}$ and also the fact that the two-body scattering length $a_{\mathrm{mm}}$ changes sign at the instability onset detuning $\bar{\nu}_{c,-}$ in Fig. 2(a).
  2. 2.
@@
    By contrast, in the narrow resonance case, the lower boundary detuning $\bar{\nu}_{c,-}$ of the unstable region in Fig. 2(b) shows a strong dependence on density. This lower boundary detuning is also the point at which the density-dependent many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$, rather than the two-body scattering length $a_{\mathrm{mm}}$, changes sign, as seen from Fig. 2(d).
  3. 3.
@@
    Interestingly, unlike in the wide resonance case, the boundary detuning $\bar{\nu}_{c,-}$ in Fig. 2(b) at moderate and high densities nearly coincides with the detuning $\bar{\nu}_{c}(n)$, which is associated with a quantum critical point (QCP) Radzihovsky et al. (2004); Romans et al. (2004) that separates the MSF from the ASF phase. This implies that proximity to the QCP plays a crucial role in the instability of these molecular superfluids.
@@
To quantitatively characterize the observations above, we derive the following three limiting forms for the scattering lengths $a_{\mathrm{mm}}$ and $a_{\mathrm{mm}}^{\mathrm{eff}}$. These formulae are appropriate in the interesting detuning regions where the compressibility changes its sign sup and they are applicable only for the magnetic field below the resonance $B<B_{0}$.
@@
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}=a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\mathrm{wide}}}{B_{0}-B}\big{)}^{2}\big{]}$ (2a)
where $\text{where }\qquad\Delta_{B}^{\mathrm{wide}}$ $=\frac{\pi}{\sqrt{6}}\sqrt{\frac{\hbar^{2}}{m_{1}(\bar{a}-a_{\mathrm{bg}})^{2}\Delta\mu_{m}}\Delta B}$ (2b)
is an effective molecular resonance width. Here, $a_{\mathrm{mm}}^{\mathrm{bg}}$, which is chosen to be positive  [^3] In this paper, we consider only the case where both $a_{\mathrm{bg}}$ and $a_{\mathrm{mm}}^{\mathrm{bg}}$ are positive., $m_{1}$, and $\Delta\mu_{m}$ are the molecular $s$-wave background scattering length, the atomic mass, and the magnetic moment difference between the open and closed channels, respectively. As in the literature  Lange et al. (2009); Chin et al. (2010) $\bar{a}$ in Eq. (2b) can be taken as $0.96R_{\mathrm{vdW}}$ where $R_{\mathrm{vdW}}$ is the van der Waals length.
@@
$a_{\mathrm{mm}}$ $=a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\big{)}^{\frac{3}{2}}\big{]}$ (3a)
where $\text{where }\quad\Delta_{B}^{\text{narrow}}$ $=\big{(}\frac{a_{\mathrm{bg}}}{a_{\mathrm{mm}}^{\mathrm{bg}}}\big{)}^{\frac{2}{3}}\big{(}\frac{m_{1}a_{\mathrm{bg}}^{2}\Delta\mu_{m}}{\hbar^{2}}\big{)}^{\frac{1}{3}}\;(\Delta B)^{\frac{4}{3}}$ (3b)
is the counterpart molecular resonance width. The scattering process underlying the resonant term in Eq. (3a)  [^4] Note that the whole contribution of the resonance term in Eq. (3a) does not depend on $a_{\mathrm{mm}}^{\mathrm{bg}}$, since the overall prefactor of $a_{\mathrm{mm}}^{\mathrm{bg}}$ is cancelled out by a factor of $1/a_{\mathrm{mm}}^{\mathrm{bg}}$ from $(\Delta_{B}^{\mathrm{narrow}})^{2}$. was illustrated in Fig. 1(b), which involves the exchange of two bosonic atom constituents between molecules. Importantly, the negative sign arises from Bose statistics. It can be shown that for a fermionic Feshbach resonance, $a_{\mathrm{mm}}$ takes a similar form to Eq. (3a), but with the negative sign in front of the resonant term replaced by a positive sign (See Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011)  [^5] On the other hand, in the fermionic studies of Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011), the effect of the molecular background interaction $a_{\mathrm{mm}}^{\mathrm{bg}}$ on $a_{\mathrm{mm}}$ was not considered. For additional discussions, see Ref. sup .).
@@
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\big{)}^{\frac{3}{2}}f(B,n)\big{]}$ (4a)
with $\text{with }\;f(B,n)$ $=1+\frac{2\sqrt{2}}{\pi}\ln\big{(}\frac{\bar{\nu}^{2}}{\bar{\nu}^{2}-\bar{\nu}_{c}^{2}(n)}\big{)},$ (4b)
where $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$ is the detuning and $\bar{\nu}_{c}(n)\approx-2\bar{\alpha}\sqrt{n}$ is its corresponding QCP value, with $\bar{\alpha}$ representing the renormalized Feshbach coupling strength defined in Eq. (6) below. Eq. (4) is applicable for the detuning less than the QCP value, $\bar{\nu}<\bar{\nu}_{c}<0$. At zero density, the QCP coincides with the resonance $\bar{\nu}_{c}=0$ and Eq. (4a) reduces to Eq. (3) for the two-body scattering length $a_{\mathrm{mm}}$.
Qualitatively one can understand why many-body physics plays an important role for the molecular condensate stability, as the main fluctuations that destabilize the MSF phase in Fig. 2 are atomic Bogoliubov excitations arising from the presence of the ground-state molecular condensate. These excitations exhibit a significant energy gap and behave like free atoms when the detuning is large and negative, as is relevant in a wide resonance. However, they become soft with a phonon-like, low-energy dispersion as the detuning approaches the QCP. For a sufficiently narrow resonance at a given density, it is this softness of the Bogoliubov excitations, rather than proximity to the two-body resonance point, that drives the instability.
@@
Variational analysis — To arrive at all of these results, we use a variational ground state analysis that incorporates both atomic and molecular condensates, as well as depletion and atomic Cooper-pairing contributions. Some aspects of these stability issues due to the presence of Cooper pairing of bosons have been discussed in prior work Evans and Imry (1969); Nozieres and Saint James (1982); Jeon et al. (2002); Koetsier et al. (2009); Lee and Lee (2010); Basu and Mueller (2008). There the emphasis was primarily restricted to single-channel models.
@@
$r_{*}$ $\equiv 4\pi\hbar^{4}/m_{1}^{2}\bar{\alpha}^{2}=\hbar^{2}/a_{\mathrm{bg}}m_{1}\Delta\mu_{m}\Delta B,$ (6)
which allows us to classify the resonance width quantitatively. If the ratio $w_{\mathrm{res}}\equiv a_{\mathrm{bg}}/r_{*}$ is much smaller than unity $w_{\mathrm{res}}\ll 1$, the resonance is viewed as narrow; otherwise, it is classified as wide. For the narrow and wide resonances used in this paper, we choose $w_{\mathrm{res}}\approx 0.006$ Wang et al. (2024); sup and $w_{\mathrm{res}}\sim 1$, respectively   [^6] We note that using an even wider resonance does not change our conclusions. We use the moderately wide resonance to make our numerical calculation easier..

### v2

@@ Abstract
Understanding and controlling interactions of ultracold molecules is a cornerstone of quantum chemistry. While the laboratory creation of degenerate molecular gases comprised of bosonic atoms has unlocked powerful new platforms for quantum simulation, progress is limited by the absence of a robust theoretical framework for characterizing inter-molecular interactions. This is in stark contrast to the situation for Fermi gases. In this Letter, we present such a framework providing universal expressions for these molecular scattering lengths as functions of experimentally measurable quantities. Our discoveries are crucial for understanding molecular condensate formation. Calculations of the compressibility reveal that a sign change in such molecular scattering lengths is directly correlated with the instability of these condensates. These results offer fresh insight with broad applications for atomic, molecular, and condensed matter physics, as well as quantum chemistry.
Introduction — Molecular condensates of bosonic atoms represent a key frontier in physics, offering a path to discover new states of matter and quantum phase transitions Radzihovsky et al. (2004); Romans et al. (2004); Radzihovsky et al. (2008); Radzihovsky and Choi (2009). The use of Feshbach resonances  Chin et al. (2010), particularly through magnetoassociation, has revolutionized our ability to create stable diatomic molecules from ultracold atomic pairs. This has allowed for the preparation of equilibrium molecular condensates Zhang et al. (2021, 2023).
Despite this progress, bosonic systems face significant challenges. Condensates become unstable near a resonance, as they experience particle loss and heating due to three-body recombination  Fedichev et al. (1996); Bedaque et al. (2000); Weber et al. (2003); Fletcher et al. (2013). Feshbach interactions can add to this condensate destabilization when they introduce an attractive force between molecules. Such instabilities contrast sharply with two-component Fermi gases, where the Pauli exclusion principle  Petrov et al. (2004, 2005) allows for the robust formation of stable molecular condensates. The stability in fermionic systems has enabled an extensive exploration of the Bardeen-Cooper-Schrieffer (BCS) to Bose-Einstein condensation (BEC) crossover  Zwerger (2012); Regal et al. (2004); Zwierlein et al. (2004); Chen et al. (2005), a level of understanding that has yet to be replicated with bosons. Here, BCS-BEC crossover refers to a superfluid undergoing a smooth evolution, with increasing interaction strength, from large, overlapping Cooper pairs to tightly bound diatomic molecules.
In this Letter, we address this critical issue of bosonic molecular condensate instability by examining the molecule-molecule scattering length. Our approach is inspired by atomic condensates, where the two-body $s$-wave scattering length is typically described by the following equation:
$a_{s}$ $=a_{\mathrm{bg}}\bigg(1-\frac{\Delta B}{B-B_{0}}\bigg),$ (1)
Here, $B_{0}$ is the resonance value of the magnetic field $B$, $\Delta B$ is the resonance width, and $a_{\mathrm{bg}}$ is the atomic $s$-wave background scattering length. The leading-order contribution to $a_{s}$ due to the Feshbach coupling can be visualized as in Fig. 1(a)  [^1] While higher-order contributions, involving repeated processes from Fig. 1(a) renormalize $B_{0}$ they don’t change the resonant structure of $a_{s}$..
@@
Figure 1: Scattering due to atomic Feshbach coupling: (a) Leading-order scattering process contributing to the resonant term in the atomic scattering length, $a_{s}$ [Eq. (1)]. Two incoming atoms (single blue circles) temporarily combine into a molecule (two attached circles) via Feshbach coupling, and then dissociate into two outgoing free atoms. The orange region represents virtual processes. (b) Corresponding leading-order contribution to the molecular scattering length, $a_{\mathrm{mm}}$, near the resonance. Here, two molecules approach each other and temporarily break up into four free atoms, which then propagate, interact, and recombine into two molecules. Similar to panel (a), this process universally depends on the Feshbach coupling but is of higher order in the coupling strength.
It is well known that atomic condensates collapse when $a_{s}$ becomes negative. This typically happens when the magnetic field $B$ is above the resonance, specifically within the range $B_{0}<B<B_{0}+\Delta B$ (assuming a positive atomic background scattering length, $a_{\mathrm{bg}}>0$). For concreteness, here we assume $\Delta B>0$, so that the bare molecules’ energy drops below the atom continuum when the magnetic field is below resonance ($B<B_{0}$)   [^2] The formulas derived here can be easily generalized to the case of negative $\Delta B$. What has not been widely appreciated is that Feshbach coupling can also modify the scattering properties of molecules, which can lead to the instability of a molecular Bose-Einstein condensate. Understanding this important effect is central to our investigation.
In this Letter, we demonstrate how Feshbach resonances dramatically alter the intermolecular scattering length and consequently the behavior of bosonic molecular condensates, mirroring their profound effect in atomic systems. Crucially, this tunability can drive molecular condensates into instability when this scattering length (called $a_{\mathrm{mm}}$ at zero and $a_{\mathrm{mm}}^{\mathrm{eff}}$ at finite density) becomes negative. Notable is a rather pronounced density dependence found in $a_{\mathrm{mm}}^{\mathrm{eff}}$ for narrow resonances, which arises from many-body effects and introduces additional complexity; only in the limit of very wide Feshbach resonances does $a_{\mathrm{mm}}^{\mathrm{eff}}$ simplify, converging with the two-body scattering length $a_{\mathrm{mm}}$.
The accurate determination of both the fundamental molecular scattering length ($a_{\mathrm{mm}}$) and the density-dependent effective scattering length ($a_{\mathrm{mm}}^{\mathrm{eff}}$) presents a significant theoretical challenge. As Figure 1(b) illustrates, even a leading-order intermolecular scattering process can be considerably more complex as it involves higher order in Feshbach coupling strength than what is seen in simpler atomic systems [Figure 1(a)]. Additionally, we also need to precisely account for background interactions between atoms and molecules.
Summary of main results — Our theoretical framework employs a two-channel variational wavefunction treatment to systematically analyze the system’s compressibility and, from it, deduce the crucial scattering lengths. We begin by presenting general insights into the zero-temperature stability phase diagrams presented in Figure 2, which are determined through numerical calculations of the compressibility $\kappa=dn/d\mu$.
@@
Figures 2(a,b) show the zero-temperature stability phase diagrams for both a wide and narrow resonance, in terms of density $n$ on the vertical axis and detuning $\bar{\nu}$ on the horizontal axis. Both panels indicate a stable molecular superfluid (MSF) phase at large negative detuning, a stable atomic superfluid (ASF) phase at large positive detuning, and an unstable region in between, where the compressibility becomes negative $\kappa<0$. The boundaries of this unstable region are marked by the detunings $\bar{\nu}_{c,+}$ and $\bar{\nu}_{c,-}$. In Fig. 2(c,d), we plot both our calculated two-body scattering length, $a_{\mathrm{mm}}$, and its many body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$.
@@
Note a striking similarity between panels (a) and (b) of Fig. 2 on the atomic side. For both narrow and wide resonances, the phase boundary of the unstable region is essentially dictated by the sign change of the two-body atomic scattering length, $a_{s}$. This means that in both scenarios, the upper boundary detuning of the unstable regime, $\bar{\nu}_{c,+}$ remains nearly independent of density.
@@
In contrast, the behavior of the compressibility on the molecular side of the phase diagrams in Figs. 2(a,b) reveals a significant difference between wide and narrow resonances. We summarize our key observations:
@@
(1) Wide resonance case with density-independent stability. In the case of wide resonances, the entire unstable region shows very little dependence on density. This is a significant finding because it means we can roughly characterize the molecular condensate’s stability using just a two-body scattering length, similarly to how we treat atomic condensates. Figure 2(c) confirms this, showing that $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ is nearly equivalent to $a_{\mathrm{mm}}$, and importantly, $a_{\mathrm{mm}}$ itself changes sign at the instability onset detuning, $\bar{\nu}_{c,-}$, in Fig. 2(a).
@@
(2) Narrow resonance case with strong density dependence. By contrast, for narrow resonances, the lower boundary detuning, $\bar{\nu}_{c,-}$, of the unstable region in Fig. 2(b) is highly dependent on density. Here, it is the density-dependent many-body scattering length, $a_{\mathrm{mm}}^{\mathrm{eff}}$, not the two-body $a_{\mathrm{mm}}$, that changes sign at this instability boundary, as clearly shown in Fig. 2(d).
@@
(3) Proximity to quantum critical point (QCP). What is particularly interesting is that for narrow resonances, at moderate to high densities, the boundary detuning $\bar{\nu}_{c,-}(n)$ in Fig. 2(b) nearly aligns with the detuning $\bar{\nu}_{c}(n)$. This detuning $\bar{\nu}_{c}(n)$ is associated with a QCP that separates the MSF from the ASF phase. This strong correlation suggests that close proximity to the QCP is a crucial factor in the instability of these narrow resonance molecular superfluids.
Figure 2: Phase diagrams and molecular scattering lengths. Ground-state stability phase diagrams for (a) wide and (b) narrow resonances, showing compressibility $\kappa$ [normalized by $\kappa_{\mathrm{bg}}=m_{1}/(4\pi\hbar^{2}a_{\mathrm{bg}})$] as a function of atom number density $n$ and of detuning $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$ (normalized by resonance width $\Delta_{\bar{\nu}}=\Delta\mu_{m}\Delta B$). Orange indicates stable regions ($\kappa>0$), and blue indicates unstable regions. The atomic condensate is present only in the atomic superfluid (ASF) phase, while the molecular condensate is present in both ASF and molecular superfluid (MSF) phases. Red dashed lines denote the quantum critical point $\bar{\nu}_{c}(n)$ separating ASF from MSF, while $\bar{\nu}_{c,-}$ and $\bar{\nu}_{c,+}$ mark boundaries between unstable and stable regions. (c),(d) Corresponding two-body molecular scattering length $a_{\mathrm{mm}}$, and its many-body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ for $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$. Parameters for the narrow resonance [panels (b) and (d)] are from the 133Cs resonance at $B_{0}=19.849$ G Zhang et al. (2021, 2023); Wang et al. (2024). For the wide resonance [panels (a) and (c)], $\Delta_{\bar{\nu}}$ is increased by $10^{2}$ times relative to panels (b) and (d).
@@
To quantitatively confirm the observations above, we present three limiting forms for the scattering lengths $a_{\mathrm{mm}}$ and $a_{\mathrm{mm}}^{\mathrm{eff}}$. These formulas are particularly relevant in the interesting detuning regions where a stable MSF exists and where the compressibility changes its sign sup . It is important to note that they are applicable only for the magnetic field below the resonance $B<B_{0}$.
@@
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}=a_{\mathrm{mm}}^{\mathrm{bg}}\bigg[1-\bigg(\frac{\Delta_{B}^{\mathrm{wide}}}{B_{0}-B}\bigg)^{2}\bigg]$ (2a)
$\text{where }\qquad\Delta_{B}^{\mathrm{wide}}$ $=\frac{\pi}{\sqrt{6}}\sqrt{\frac{\hbar^{2}}{m_{1}(\bar{a}-a_{\mathrm{bg}})^{2}\Delta\mu_{m}}\Delta B}$ (2b)
is an effective molecular resonance width. Here, $a_{\mathrm{mm}}^{\mathrm{bg}}$, which is chosen to be positive  [^3] In this Letter, we consider only the case where both $a_{\mathrm{bg}}$ and $a_{\mathrm{mm}}^{\mathrm{bg}}$ are positive., $m_{1}$, and $\Delta\mu_{m}$ are the molecular $s$-wave background scattering length, the atomic mass, and the magnetic moment difference between the open and closed channels, respectively. As in the literature Lange et al. (2009); Chin et al. (2010), $\bar{a}$ in Eq. (2b) can be taken as $0.96R_{\mathrm{vdW}}$, where $R_{\mathrm{vdW}}$ is the van der Waals length.
@@
$a_{\mathrm{mm}}$ $=a_{\mathrm{mm}}^{\mathrm{bg}}\bigg[1-\bigg(\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\bigg)^{\frac{3}{2}}\bigg]$ (3a)
$\text{where }\quad\Delta_{B}^{\text{narrow}}$ $=\bigg(\frac{a_{\mathrm{bg}}}{a_{\mathrm{mm}}^{\mathrm{bg}}}\bigg)^{\frac{2}{3}}\bigg(\frac{m_{1}a_{\mathrm{bg}}^{2}\Delta\mu_{m}}{\hbar^{2}}\bigg)^{\frac{1}{3}}\;(\Delta B)^{\frac{4}{3}}$ (3b)
is the counterpart molecular resonance width. The scattering process underlying the resonant term in Eq. (3a)  [^4] Note that the whole contribution of the resonance term in Eq. (3a) does not depend on $a_{\mathrm{mm}}^{\mathrm{bg}}$, since the overall prefactor of $a_{\mathrm{mm}}^{\mathrm{bg}}$ is cancelled out by a factor of $1/a_{\mathrm{mm}}^{\mathrm{bg}}$ from $(\Delta_{B}^{\mathrm{narrow}})^{2}$. is illustrated in Fig. 1(b), which involves the exchange of two bosonic atom constituents between molecules. Importantly, the negative sign in front of this term is fundamentally connected to Bose statistics. It can be shown that for a fermionic Feshbach resonance, $a_{\mathrm{mm}}$ takes a similar form to Eq. (3a), but with the negative sign in front of the resonant term replaced by a positive sign (see Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011)  [^5] On the other hand, in the fermionic studies of Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011), the effect of the molecular background interaction $a_{\mathrm{mm}}^{\mathrm{bg}}$ on $a_{\mathrm{mm}}$ was not considered. For additional discussions, see Ref. sup .).
@@
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}^{\mathrm{bg}}\bigg[1-\bigg(\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\bigg)^{\frac{3}{2}}f(B,n)\bigg]$ (4a)
$\text{with }\;f(B,n)$ $=1+\frac{2\sqrt{2}}{\pi}\ln\bigg(\frac{\bar{\nu}^{2}}{\bar{\nu}^{2}-\bar{\nu}_{c}^{2}(n)}\bigg),$ (4b)
where $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$ is the detuning and $\bar{\nu}_{c}(n)\approx-2\bar{\alpha}\sqrt{n}$ is its corresponding QCP value, with $\bar{\alpha}$ representing the renormalized Feshbach coupling strength defined in Eq. (6). Equation (4) is applicable for the detuning less than the QCP value, $\bar{\nu}<\bar{\nu}_{c}<0$. At zero density, the QCP coincides with the resonance $\bar{\nu}_{c}=0$ and Eq. (4a) reduces to Eq. (3) for the two-body scattering length $a_{\mathrm{mm}}$.
Qualitatively, one can understand why many-body physics plays an important role in the molecular condensate stability, as the main fluctuations that destabilize the MSF phase in Fig. 2 are atomic Bogoliubov excitations arising from the presence of the ground-state molecular condensate. These excitations exhibit a significant energy gap and behave like free atoms when the detuning is large and negative, as is relevant in a wide resonance. However, they become soft with a phonon-like, low-energy dispersion as the detuning approaches the QCP. For a sufficiently narrow resonance at a given density, it is this softness of the Bogoliubov excitations, rather than proximity to the two-body resonance point, that drives the instability.
@@
Variational analysis — To arrive at all of these results, we use a variational ground state analysis that incorporates both atomic and molecular condensates, as well as depletion and atomic Cooper-pairing contributions. Some aspects of these stability issues due to the presence of Cooper pairing of bosons have been discussed in prior work Evans and Imry (1969); Nozieres and Saint James (1982); Jeon et al. (2002); Koetsier et al. (2009); Lee and Lee (2010); Basu and Mueller (2008). There, the emphasis was primarily restricted to single-channel models.
@@
$r_{*}$ $\equiv 4\pi\hbar^{4}/m_{1}^{2}\bar{\alpha}^{2}=\hbar^{2}/(a_{\mathrm{bg}}m_{1}\Delta\mu_{m}\Delta B),$ (6)
which allows us to classify the resonance width quantitatively. If the ratio $w_{\mathrm{res}}\equiv a_{\mathrm{bg}}/r_{*}$ is much smaller than unity $w_{\mathrm{res}}\ll 1$, the resonance is viewed as narrow; otherwise, it is classified as wide. For the narrow and wide resonances used in this Letter, we choose $w_{\mathrm{res}}\approx 0.006$ Wang et al. (2024); sup and $w_{\mathrm{res}}\sim 1$, respectively   [^6] We note that using an even wider resonance does not change our conclusions. We use the moderately wide resonance to make our numerical calculation easier..